OpenStudy (anonymous):

A radio active isotope decays according to the exponential decay equation where t is in days. Round to the thousandths place. For the half life: The half life is the solution (t) of the equation : a/2=ae^−4.457t

OpenStudy (danjs):

\[A = A _{0}*e^{r*t}\] you want A to be half the original amount A sub zero. that is the equation they gave you

OpenStudy (danjs):

notice now, both sides have an a, divide both sides by a to eliminate that

OpenStudy (danjs):

\[\frac{ 1 }{ 2 }=e^{-4.457*t}\]

OpenStudy (anonymous):

now you do loge^1/2=-4.457?

OpenStudy (danjs):

log base e is the natural log, ln you can do that to both sides

OpenStudy (danjs):

\[\ln(\frac{ 1 }{ 2 })=\ln(e^{-4.457*t})\]

OpenStudy (anonymous):

0.6931=-1.4944?

OpenStudy (danjs):

hmm, do you remember the log exponent rule.. \[\log_{b}(x^a) = a *\log_{b} (x) \]

OpenStudy (danjs):

and also, \[\log_{b}(b) = 1 \]

OpenStudy (danjs):

so on the right, pull down the exponent into the front of the log... and replace the natural log of e with 1.

OpenStudy (anonymous):

im not good at this at all :(

OpenStudy (danjs):

\[\ln(\frac{ 1 }{ 2 })=\ln(e^{-4.457*t}) = -4.457t * \ln(e)\]

OpenStudy (danjs):

just have to remember the rules for logs

OpenStudy (danjs):

ln(e) = 1, so ln(1/2) = -4.457*t just have to solve for t, with a calculator

OpenStudy (anonymous):

1.5550?

OpenStudy (danjs):

yep, just round to the thousandths place, 1.555 days

OpenStudy (anonymous):

Could you help me with one more please?

OpenStudy (anonymous):

e^x+1=1.038?

OpenStudy (danjs):

is e^(x+1) all exponent, or just x

OpenStudy (anonymous):

x+1

OpenStudy (danjs):

similar to last prob.. take the ln of both sides ln[e^(x+1)] = ln(1.038)

OpenStudy (danjs):

exponent rule brings (x+1) to the front (x+1) * ln(e) = ln(1.038)

OpenStudy (danjs):

natural log of e is 1, so x + 1 = ln(1.038) x = ln(1.038) + 1

OpenStudy (anonymous):

1.0372?

OpenStudy (danjs):

yep

OpenStudy (anonymous):

then rounded to the nearest thousandth it would just be 1.037?

OpenStudy (danjs):

yes

OpenStudy (anonymous):

thank you!!!